# DB3 Team’s Solution For Meta KDD Cup’ 25

### v1

@@ Abstract.
copyright: acmlicensedjournalyear: 2025conference: 2025 KDD Cup Workshop; isbn: ;$c$s: Computing methodologies Natural language generation
@@ 1. Introduction
Recent advancements in artificial intelligence have highlighted the importance of multi-modal reasoning, especially in scenarios where systems must interpret and respond to both visual and textual information. The Meta CRAG-MM Challenge 2025, hosted on AIcrowd, aims to accelerate progress in this area by introducing a novel benchmark, called CRAG-MM, that encompasses both single-turn and multi-turn visual question answering tasks. Unlike traditional VQA datasets, CRAG-MM is designed to evaluate factual question answering in multi-modal contexts, integrating images with complex, multi-turn conversational flows.
@@
Train/Validation Split. To perform checkpoint selection and some other techniques, we have to split the training and validation sets. As the query types and domains vary, we also split the train/validation set proportionally to make sure each question type and domain appears equally in the train and validation splits.
@@ 3. The Retrieval Component of the Solution
In Task 1, an image-indexed knowledge graph is provided as the information source. In Task 1, an image-indexed knowledge graph is provided as the information source. Specifically, an image is linked with a CLIP (Radford et al., 2021) index and is associated with a structured segment of text, which serves as a description of the entity represented by the image. The information useful for answering the question is mostly based on the information in the structured text; therefore, the task involves identifying the correct entity the query is about in the knowledge graph.
@@
Specifically, we sample the top 5 candidates for each query using the grounding and retrieving through image pipeline. We verify each candidate by using them to retrieve content to answer the query and judging through a strong VLM, e.g., GPT-4o. A candidate is viewed as a related item if the retrieved item helps the base VLM answer the query correctly, and the strong VLM judges the entity in the index image to be the same as the entity in the grounded query image. Such a candidate is labeled as relevant, and the base VLM is trained to classify the relevance.
@@
1Given an image and a query about it, your task is to extract the entitys name the query is about.
2If the entity is about a plant, you should output the full botanical name.
3....
4If the enityt is about a vehicle, you should output the vehicles brand and model.
5Query: {}
6You should output the entitys name directly.
@@
1###Task You are an expert at converting visual questions into effective search queries.
2Your goal is to create a comprehensive search query that will help find the most relevant information.
3For each image-based question, you must create a search query that combines:
@@ 63. Potential answer terms or relevant context
7For example:
8- If asking about a logo: include company name, industry, and visual description
9- If asking about an object: include its appearance, category, and possible brands/models
10- If asking about an event/scene: include location hints, activities, and time period clues
11’’’
12Image:{}
@@
We tune the Llama 3.2-VL base model for better rewrite performance. We select both the cheated and non-cheated rewrite queries as the training sample for Llama 3.2-VL. The results are also presented in Table 1. As we can see, regardless of using the cheated or non-cheated version, the boost of potential performance is 4̃%. The performance is close to the GPT-4o rewrite performance ( 3% down); therefore, the result is rather satisfying. We choose one of the checkpoints in the Llama-distill checkpoints as the final rewrite module we use in the submitted version.
@@
1###Task You are an expert at converting visual questions into effective search queries.
2The current query is a part of multi-turn conversation. You should use the history conversation to make sure what the current query is about.
3Your goal is to create a comprehensive search query that will help find the most relevant information for the currecnt query.
4For each image-based question, you must create a search query that combines:
@@ 73. Potential answer terms or relevant context
8For example:
9- If asking about a logo: include company name, industry, and visual description
10- If asking about an object: include its appearance, category, and possible brands/models
11- If asking about an event/scene: include location hints, activities, and time period clues
12Query: {}
13Context: {}
@@ 4.2. Refusal Training Pipeline
We obtain different checkpoints using different queries for further selection and ensemble.
@@ 7. Acknowledgement
Table 2. LoRA and FineTuning Hyperparameters.
Name Value
LoRA_alpha 16
LoRA_dropout 0.1
LoRA_r 8
target_modules
[”k_proj”, ”q_proj”, ”v_proj”,
”up_proj”, ”down_proj”, ”gate_proj”]
bias ”none”
4-bit True
max_seq_length 2048/4096
per_device_train_batch_size 1
gradient_accumulation_steps 4
optim ”adamw_hf”
learning_rate 2e-4
max_grad_norm 0.3
scheduler ”cosine”, warm_up_ratio=0.1

### v2

@@ Abstract.
copyright: acmlicensedjournalyear: 2025conference: 2025 KDD Cup Workshop; $c$s: Computing methodologies Natural language generation
@@ 1. Introduction
Recent advancements in artificial intelligence have highlighted the importance of multi-modal reasoning, especially in scenarios where systems must interpret and respond to both visual and textual information. The Meta CRAG-MM Challenge 2025 (Wang et al., 2025) hosted on AIcrowd, aims to accelerate progress in this area by introducing a novel benchmark, called CRAG-MM, that encompasses both single-turn and multi-turn visual question answering tasks. Unlike traditional VQA datasets, CRAG-MM is designed to evaluate factual question answering in multi-modal contexts, integrating images with complex, multi-turn conversational flows.
@@
Train/Validation Split. To perform checkpoint selection and other validation-based techniques, we split the original training data into distinct training and validation subsets. Given the diverse query types and domains present in the dataset, it is crucial to ensure that both splits are representative of the overall data distribution. To achieve this, we employ a stratified splitting strategy based on question types and domains, ensuring that each category appears proportionally in both the training and validation sets. Specifically, we allocate about 3,000 of the samples to the training set and 800 to the validation set, maintaining an 8:2 ratio across all types and domains. This approach mitigates the risk of overfitting to specific data characteristics and supports more reliable model evaluation and checkpoint selection.
@@ 3. The Retrieval Component of the Solution
In Task 1, an image-indexed knowledge graph is provided as the information source. Specifically, an image is linked with a CLIP (Radford et al., 2021) index and is associated with a structured segment of text, which serves as a description of the entity represented by the image. The information useful for answering the question is mostly based on the information in the structured text; therefore, the task involves identifying the correct entity the query is about in the knowledge graph.
@@
We sample the top 5 candidate items for each query in the training set using our image-based grounding and retrieval pipeline. Each candidate is verified by a powerful VLM (e.g., GPT-4o) to determine whether the entity in the index image matches the entity grounded in the query image. A candidate is labeled as relevant only if this condition is satisfied. This process resulted in a distillation dataset of approximately 15,000 candidate-query pairs, within which roughly 500 positive samples were identified for Task 1. The base VLM is subsequently trained on these data to classify the relevance of candidate items. To address class imbalance, we employ negative sampling during training to construct mini-batches with an approximately 1:1 ratio of positive to negative samples. This balanced sampling strategy ensures that the model receives representative supervisory signals from both relevant and irrelevant instances at each optimization step, thereby mitigating bias toward the majority class.
@@
1Image:{}
2Given an image and a query about it, your task is to extract the entitys name the query is about.
3If the entity is an animal, output its scientific name.
4If the entity is a plant, output its full botanical name.
5If the entity is food, output the specific dish or ingredient name.
6If the entity is a location or landmark, output its common name.
7If the entity is a mathematical concept or symbol, output its formal name.
8If the enityt is about a vehicle, you should output the vehicles brand and model.
9For other entities, output the most specific and commonly recognized name.
10Query: {}
11You should output the entitys name directly.
@@
1Image:{}
2###Task
3You are an expert at converting visual questions into effective search queries.
4Your goal is to create a comprehensive search query that will help find the most relevant information.
5For each image-based question, you must create a search query that combines:
@@ 83. Potential answer terms or relevant context
9For example:
10- If asking about a logo: include company name, industry, and visual description
11- If asking about an object: include its appearance, category, and possible brands/models
12- If asking about an event/scene: include location hints, activities, and time period clues
@@
14Answer: {} //When cheating, it will provide the answer.
@@
We tune the Llama 3.2-VL base model to enhance its query rewriting capability. For the training data, we utilize both cheated and non-cheated rewrite queries, but critically, we only employ those queries for which GPT-4 generated a correctresponse as the ground-truth label to supervise the fine-tuning process of Llama 3.2-VL. This ensures that the model learns from accurate and reliable demonstrations, improving its ability to generate effective query rewrites. The results are also presented in Table 1. As we can see, regardless of using the cheated or non-cheated version, the boost of potential performance is 4̃%. The performance is close to the GPT-4o rewrite performance ( 3% down); therefore, the result is rather satisfying. We choose one of the checkpoints in the Llama-distill checkpoints as the final rewrite module we use in the submitted version.
@@
1Image:{}
2###Task
3You are an expert at converting visual questions into effective search queries.
4The current query is a part of multi-turn conversation. You should use the history conversation to make sure what the current query is about.
5Your goal is to create a comprehensive search query that will help find the most relevant information for the currecnt query.
6For each image-based question, you must create a search query that combines:
@@ 93. Potential answer terms or relevant context
10For example:
11- If asking about a logo: include company name, industry, and visual description
12- If asking about an object: include its appearance, category, and possible brands/models
13- If asking about an event/scene: include location hints, activities, and time period clues
14Query: {}
15Context: {}
@@ 4.2. Refusal Training Pipeline
We conduct comprehensive experiments with all three approaches— SFT, DPO, and GRPO. The results indicate that the SFT approach performs poorly on this task, achieving a negative score. For Task 1, both DPO and GRPO yield comparable final scores of approximately 5%. However, we observe notable behavioral differences: models trained with DPO tend to be more conservative, exhibiting a high missing rate (70%-90%) as they frequently respond with ”I don‘t know.”. In contrast, models trained with GRPO adopt a more proactive response strategy, demonstrating a lower missing rate (60%-80%) and showing a greater willingness to attempt answers. Additionally, each method displays distinct strengths across different domains. Given these complementary characteristics, we subsequently explore an ensemble strategy to leverage the advantages of both methods.